OpenStudy (anonymous):

Hey can someone help me with some algebra? I just need some help on this: Kyra observed that 85% of the students at a library liked reading mystery stories. If 30 students at the library did not like reading mystery stories, the number of students at the library who liked mystery stories was _____. (only put numeric values, no other symbols)

OpenStudy (anonymous):

30 students did NOT like reading mystery stories. 100 - 85 = 15. So 15% of the students did not like reading mystery stories.

OpenStudy (anonymous):

170.

OpenStudy (anonymous):

85 / 15 = 17/3 30 * 17/3 = 170.

OpenStudy (anonymous):

So 85% = 170. And 30 = 15%

OpenStudy (anonymous):

Now you understand how to figure out the problem, correct? @SectorZ423

OpenStudy (anonymous):

let the number of students be x (85/100)*x=0.85x remaining students=x-0.85x=0.15x 0.15x=30 x=200

OpenStudy (anonymous):

Correct now (just because) how many students in total?

OpenStudy (anonymous):

200

OpenStudy (anonymous):

170 + 30 = 200. 85% + 15% = 100%.

OpenStudy (anonymous):

thanks even though that wasnt part of the answer I just wanted to know

OpenStudy (anonymous):

so the number of students who liked mystery stories=200-30=170

OpenStudy (anonymous):

Exactly